Wednesday, April 5, 2017

From HallCWiki
Jump to: navigation, search

Meeting Information

  • Location CC L210A, 11 am
  • DocDB

Blue Jeans Connection Information

  1. Enter Meeting ID: 895800575
  2. Press #

Agenda

  1. Status update from detector owners, calibrations, efficiencies, etc.
    • Review systems owners and update information if needed
  2. Address system specific points of interest for the Fall 2017 run

Minutes

General

  • All groups are requested to update documentation on the wiki link
    • We will use these as references for shift crew later.
    • See links under System Owners
    • Use the HMS Drift Chambers entries as an example
    • Links should point to documents in the DocDB
    • Don't just upload to the wiki
  • Please ensure any hardware changes you make get put in the HCLog
    • Even if they shouldn't matter to others, it is very useful to have the history logged *AND* a record of where people have been working

Miscellaneous

  • We will need commissioning plans for each detector in the stack, this will be the responsibility of the detector owners
  • Survey and alignment will need to be conducted this summer
  • Engineering work is required for the design of an access platform which allows one to access the detector modules in the stack
    • This work should be coordinated with the previous design of "Simona's Platform"
  • If there are EPICS variables that folks would like to be incorporated into the DAQ, they should contact Brad S.
    • All EPICS values of interest need to be sanity checked with the Spring 2017 commissioning data so that issue can be identified and resolved
    • e.g. the SHMS wire chamber currents were not being recorded during the Spring 2017 commissioning run

Trigger/DAQ

  • Steve will be playing with the HMS DAQ in order to integrate the FADC scalers into the readout
    • If the DAQ is needed for any reason, contact Steve W. to give the DAQ back
      • It is an instantaneous operation
  • The coincidence trigger still needs to be tested
    • This will be completed this summer
  • There are 4 SHMS pre-shower sums already patched to the counting house
  • The SHMS Cherenkov detectors have not yet been connected to 1190s but this will happen within the next few weeks

Beam Line

  • BCM's and BPM's status:
    • Need a wire calibration, Dave will check the software
    • There are a few more that will need to be hooked up
    • A few were saturating however this is an easy fix
    • Legacy things were working during the Spring run
    • New digital receivers were saturating
    • One more needs to be hooked up but Dave has the cables for these
  • No bulls eye scan was performed during the spring run, a plan is needed for the Fall run
  • Dave G. is checking that the harps are being re-surveyed
  • We will need cross checks on energy measurements during fall run

HMS

6 GeV Drift Chambers

  • The HMS drift chambers have been calibrated using the HMS run 303 from KPP
  • The new time-to-distance maps for each of the twelve planes as well as the t0-offsets have been up- dated in the hallc-replay parameter files hdc tracking.param (t0-offsets) and hdriftmap.param (time-to-distance maps)
  • The t0 offsets had to be re-adjusted manually in order to obtain flat drift distance histograms, as there are significant differences in the individual wire t0s, which makes the drift distances non-flat at the edges
  • Chambers will likely be stored in the ESB, the exact location is unknown

12 GeV Drift Chambers

  • New cables will have to be either bought or stolen
    • Same 1190 readout and disc cards as used with old chambers
  • The status of the new HMS chambers is not known

Hodoscopes

  • Mark J. reported that they the data from the spring run look fine
  • Simona is working on calibration

Shower

  • Calibrated okay, no problems

Heavy Gas Cherenkov

  • It looks like it should, there is a known inefficiency in the middle of chamber, this is okay

Aerogel

  • Arthur cleaned it
  • Tanja H. has spare aerogel and PMTs that could be used for replacement parts
  • Residing in EEL in Walter's shop
    • Possibly ready by fall although it is not needed
  • The new WC frame will support the aerogel
  • Plan is to repair the detector this summer with students

SHMS

Drift Chambers

  • Deb should be done with classes soon so he will be able to spend more time on the chambers soon
  • Deb has been focusing on the t0 offsets, wire by wire,
  • Decided to do by averaging connector by connector however, it will eventually be wire by wire
  • The procedure is already available in the calibration code
  • Deb will talk with Mark to get this integrated into the hallc_replay and then make the necessary modifications to hcana to read the new parameters
  • They are going to open the spare chamber sometime next week so that they can locate the gas leak

GEM

  • Work is ongoing to incorporate the GEM into the SHMS DAQ via pedestal data

Hodoscopes

  • System Owners: S. Malace
  • Simona is working on the calibration of the plastic scintillators
  • The plan is to replace all the tubes in the quartz plane
  • There is a 12 week lead time on the tubes; the order has not been placed yet, Steve W. is waiting for a purchasing code from Allison
  • Simona is working on a test setup in Hall C; this setup will be used for cosmic tests of the quartz bars once they are fitted with new PMTs AND ALSO for various tests of the new PMTs once they arrive on site
    • Steve W. is working on the DAQ for this setup
  • Simona updated the SHMS hodoscope wiki, there are two sets of slides
  • The main observation was that there are a number of tubes that have a much lower yield than expected

Shower/Pre-shower

  • Everything is okay for now according to beam data
  • They are thinking of something to do this summer, they require special trigger
  • They do not want to increase HV in order to look at cosmic data
  • There was not enough beam time in Spring to play around with calorimeter
  • While pions and electrons are visible in the calorimeter as a whole, it is not visible in the individual blocks
  • They will look at data when the quads were defocused to study blocks that were not illuminated with the hour glass focal plane distribution
  • Take home message, shielding in front of the dipole needs to be improved prior to the Fall run
  • Dave presented slides for the anode current study that he performed
  • Take home message is were were running at low luminosity and with the current draw at the level it is now, we will not be able to run this way in the fall

Aerogel Cerenkov

  • Working fine, nice distributions, nice calibrations

NGC

  • Eric and Donal will meet next Wednesday to discuss future commissioning plans

HGC

  • Ryan A. reported calibration is ongoing, tracking cuts and PID cuts between the pre-shower and NGC are being studied
  • Ryan is still taking classes however he will have more time soon
  • The calibration scripts should be easily adapted to the NGC